# Preference-driven Knowledge Distillation for Few-shot Node Classification

### v1

@@ 1 Introduction
Text-attributed graphs (TAGs yang2021graphformers ), such as citation, webpage, and product graphs feng2024taglas ; li2025hetgb , have nodes associated with text attributes. Graph neural networks (GNNs) kipf2016semi ; yun2019graph have demonstrated excellent performance and efficiency in node classification on TAGs, which are supported by high-quality labels and effective message-passing mechanisms zhu2021textgnn . However, the manual labeling of nodes is undoubtedly a tedious, expensive, and time-consuming task shi2025latent ; cao2025low . In many scenarios, only a few node labels are available. Additionally, nodes often have complex and diverse interaction relationships with each other—their local topologies are intricate—which challenge traditional GNNs with fixed message-passing mechanisms. Compared with GNNs, large language models (LLMs) exhibit impressive zero-/few-shot learning capabilities on TAGs chen2024exploring ; wang2024llms ; wu2024exploring . But the large parameter scale considerably hinders their inference efficiency he2023harnessing .
A natural idea is to blend their complementary strengths for few-shot node classification on TAGs. Knowledge distillation (KD) hinton2014distilling is a feasible solution. However, directly distilling knowledge from the LLM to GNN is impractical. Firstly, the discrepancy of decoder-only (LLMs) and encoder-only (GNNs) leads to fundamentally different characteristics in their embedding spaces chen2023label . And the huge embedding-dimension difference needs sophisticated embedding alignment and also brings high training cost hu2025large . In contrast, conducting prediction distillation from LLMs to GNNs by annotating node labels can efficiently alleviate the label scarcity and scalability dilemma pan2024distilling . The critical question is how to select the nodes for the LLM’s label annotation to effectively enhance teacher GNNs. Generally, one may use uncertainty fox1957training as a selection metric in the embedding space of GNN. However, owing to nodes’ diverse semantic and complex structural attributes (e.g., local topologies), a single GNN cannot capture the essences of nodes completely liu5084903distilling . Therefore, we investigate the embedding spaces of various-architecture GNNs to effectively mitigate cognitive limitations wang2021mulde associated with relying on a single GNN, thereby better selecting nodes for LLM’s label annotations.
@@
To this end, we propose a preference-driven knowledge distillation (PKD) framework that unites the complementary strengths of LLMs and various-architecture GNNs for few-shot node classification on TAGs. It mainly includes two modules: GNN-preference-driven Node Selector (GNS) and Node-preference-driven GNN Selector (NGS). The prerequisite of GNS is that the LLM should be able to comprehend the graph topology. Thus, we develop the graph topology aware (GTA) prompts to fine-tune the LLM, enhancing its capacity to comprehend graph topology. GNS fully exploits nodes’ prediction discrepancies among various GNNs to decide nodes whose labels are annotated by the LLM will effectively enhance teacher GNNs, facilitating knowledge distillation from the LLM to teacher GNNs. NGS selects for each node the most appropriate GNN message-passing mechanism, facilitating the tailored knowledge distillation from various teacher GNNs to the student GNN. It regards the fine-tuned LLM as the RL-based (reinforcement learning) agent, which treats all textualized node-specific attributes (including node’s semantic, structure, and prediction attributes) as state and the student GNN’s performance as reward.
Contributions
❶We introduce a preference-driven knowledge distillation (PKD) framework to synergize the complementary strengths of the LLM and various GNNs ingeniously for few-shot node classification on TAGs; ❷ we propose a GNN-preference-driven node selector, effectively determining nodes for annotation by the LLM and promoting knowledge distillation from the LLM to teacher GNNs; ❸ we propose a node-preference-driven GNN selector to tailor for each node the most appropriate message-passing mechanism, promoting knowledge distillation from teacher GNNs to the student GNN; ❹ we validate the efficacy of PKD for few-shot node classification on nine TAGs. The experiments show that it even defeats some state-of-the-art methods that use more node labels.
@@ 2.1 Graph Neural Networks
The field of graph learning has been dominated by GNNs, such as GCN kipf2016semi , GAT velivckovic2017graph , APPNP gasteiger2018predict , $\text{H}_{2}\text{GCN}$ zhu2020beyond , and GPRGNN chien2020adaptive . Recently, HoloNets koke2023holonets introduces a dual-filter mechanism with spectral response, extending spectral convolutions to directed graphs. DirGNN rossi2024edge defines the in-neighbors and out-neighbors and performs separate propagation and aggregation, improving the message passing through the incorporation of edge directionality. To deal with label scarcity, GCNII chen2020simple introduces initial residual connections and identity mapping to construct a deep GNN while EGNN satorras2021n enforces equivariance constraints for the enhancement of data efficiency and generalization. AGST ding2024toward and IceBerg li2025iceberg leverage the different self-training li2018deeper methods to effectively utilize unlabeled nodes.
@@ 2.2 Knowledge Distillation
KD is not only used for model compression, but for strengthening purposeful abilities of the student model. GFL yao2020graph extracts structural knowledge from a pre-prepared similar auxiliary graph, distilling it to the target graph for enhancing few-shot node classification performance. KDGA wu2022knowledge utilizes multiple graph augmentation strategies to make student GNN produce robust node representations after distillation. MSKD zhang2022multi mitigates the diverse classification situations requiring for different nodes by capturing multi-scale topological semantics distilled from varying layers. However, the capability of an individual teacher is inherently limited. BGNN guo2023boosting distills complementary knowledge from multiple GNN teachers sequentially and integrate it by the adaptive temperature parameter and weight boosting modules. MTAAM yang2024two distills knowledge of multiple teacher GNNs into an MLP-student, offering quick inference speed without compromising accuracy. FairGKD zhu2024devil obtains equitable and informative node representations by synergizing multiple GNN experts into a teacher. DMKD liu5084903distilling harnesses complementary knowledge from various GNNs and conducts layer-level knowledge distillation to mitigate the constraint of a single teacher. Furthermore, chen2023label is a label-free method that proposes the LLM-GNN. It uses LLMs to get high-quality annotation through active and confidence-awareness node selection, thereby circumventing the difficulty of label annotation by humans. LinguGKD hu2025large introduces a kind of ingenious contrastive learning to align the LLM’s semantic features with GNN’s structural features to achieve knowledge transfer. Most of the above knowledge distillation methods do not tailor for each node the most appropriate message-passing mechanism and underperform on few-shot node classification.
@@ 4.1 Experimental Setup
In order to assess the few-shot node classification performance of our method on TAGs, we conduct a comprehensive series of experiments across 9 real-world datasets: Cornell, Washington, Texas, Wisconsin zhu2020beyond , Amazon Ratings platonov2023critical , Ogbn-Arxiv hu2020open , Wiki CS mernyei2020wiki , Pubmed, Cora yang2016revisiting . They have various 1-hop homophily ratios yang2022incorporating and additional details of the datasets can be found in Appendix A. For the KD-baselines, we partition the nodes of each graph into training, validation, and test sets, allocating 48%, 32%, and 20%, respectively, based on the proportion division mentioned in liang2024sign . For PKD and other baselines, we randomly select 1, 3, and 5 labeled nodes per class as the initial training data and then expand the dataset to 48% of the total using the GNS module. The remaining data is randomly split into 32% for validation and 20% for testing, with the preserved indices for the baselines. This operation is repeated 5 times. We report the average test classification accuracy and standard deviation of each model with parameters that lead to the peak validation accuracy.
@@
We compare our method against the following baseline models: (i) Advanced GNNs: GCNII chen2020simple and EGNN satorras2021n ; (ii) GNNs enhanced by LLMs: LLMGNN chen2023label and GAugLLM fang2024gaugllm ; (iii) self-training for graph learning: Self-training li2018deeper , AGST ding2024toward and IceBerg li2025iceberg ; (iv) Knowledge Distillation (KD) for GNNs: KDGA wu2022knowledge , MSKD zhang2022multi , BGNN guo2023boosting , MTAAM yang2024two , and FairGKD zhu2024devil . For homophily graphs, the teacher GNNs used are: GCN kipf2016semi ($T_{1}$), GAT velivckovic2017graph ($T_{2}$), APPNP gasteiger2018predict ($T_{3}$), $\text{H}_{2}\text{GCN}$ zhu2020beyond ($T_{4}$), and the student is GCN; for heterophily graphs, the teacher GNNs employed are: DirGNN rossi2024edge ($T_{1}$), GPRGNN chien2020adaptive ($T_{2}$), HoloNets koke2023holonets ($T_{3}$), $\text{H}_{2}\text{GCN}$ ($T_{4}$) and the student is $\text{H}_{2}\text{GCN}$. The LLM used in the experiments is Llama-3.1-8B-Instruct dubey2024llama .
@@ 45.02 Amazon Ratings
To assess the effectiveness of each part in the reward function (Equation (4)), we visualize the training processes of three variants in Figure 4: (a) $R_{1}$: The reward function for teacher GNN selection depends solely on the classification accuracy of the student GNN on the expanded training data; (b) $R_{2}$: In addition to classification accuracy, the reward function also incorporates the negative cross-entropy loss ($\mathcal{L}_{CE}$); (c) $R_{3}$: Building upon $R_{2}$, the reward function also includes the negative knowledge distillation loss ($\mathcal{L}_{DL}$). As shown in Figure 4, both the three parts contribute to the improved classification performance.
@@ 4.5 Running Time
Datasets / Methods Self-training AGST IceBerg KDGA MSKD BGNN MTAAM FairGKD PKD
@@ Appendix A Detailed Description of Datasets
Cornell, Washington, Texas, Wisconsin:
@@
Cora, Pubmed and Ogbn-Arxiv:
@@ B.3 Graph Topology Aware (GTA) Prompts
The full prompts for Connectivity is presented below. When generating prompts for different datasets, we adjust certain descriptions to better align with the specific context. For example, when constructing prompts for Texas, the background description should be adapted to reflect web pages, and the relationship should be revised to hyperlinks, along with other context-specific adjustments. Similarly, for each task, the prompts must also be modified to correspond to the specific content described in Section 3.2.
@@ Appendix C Implementation Details and Time Complexity Analysis
1 $\tilde{\mathcal{D}_{L}}\leftarrow\mathcal{D}_{L}$ // Training dataset initialization ;
// LLM fine-tuning;
@@ 11      Store ($\mathcal{P}_{i},\mathbf{m}_{i},R_{i}$) to the episode history $\mathcal{F}$;
The time complexity of PKD training is mainly divided into three parts: LLM fine-tuning, GNN-preference-driven Node Selector and Node-preference-driven GNN Selector. The GNN-preference-driven Node Selector also can be divided into the annotations generation and prediction distillation.
@@
Dataset size Epoch lora_r lora_alpha optimizer learning rate GTA fine-tuning time LLM
53,617 2 4 4 AdamW $1e^{-4}$ 9h 41m 48s Llama-3.1-8B-Instruct
@@
The time complexity of Node-preference-driven GNN Selector is $\mathcal{O}(W(l^{2}d+ld^{2}+dd’+d^{\prime}a))$, where $d^{\prime}$ is the dimension of the MLP hidden layer, $a$ is the number of action categories, $W\ll l$, $a\ll d$, so the time complexity is bound by $\mathcal{O}(l^{2}d+ld^{2}+dd^{\prime})$. The training time complexity of student GNN is $\mathcal{O}((NF+M)D)$ Therefore, the overall time complexity is bound by $\mathcal{O}((L+2l)d^{2}+(d’+nL)d+2l^{2}+2D(NF+M))$.
@@
Table 13: The time costs on Cora and Ogbn-Arxiv of different LLMs.
Dataset Llama-3.1 Qwen-2.5
Parameters 8B 13B 7B 14B
Cora 0.11 GPU-hours 0.36 GPU-hours 0.05 GPU-hours 0.29 GPU-hours
Ogbn-Arxiv 1.68 GPU-hours 5.01 GPU-hours 0.94 GPU-hours 2.95 GPU-hours
Table 14: The time costs of retraining teacher GNNs on Cora, Amazon Ratings, and Ogbn-Arxiv.
Datasets Teacher GNNs Epochs Total running time (seconds)
Cora GCN, GAT, APPNP, H2GCN 300 2.4716
Obgn-Arxiv GCN, GAT, APPNP, H2GCN 300 18.2017
Amazon Ratings DirGNN, GPRGNN, HoloNets, H2GCN 300 3.1452
Table 15: Running times of PPO on Cora and Obgn-Arixv. "m" and "s" denote minute and second.
Dataset Peak memory Running time / epoch
Cora 454.62 MB 7.9s
Obgn-Arixv 1655.18 MB 44m 8.3s

### v3

@@ 1 Introduction
Text-attributed graphs (TAGs yang2021graphformers ), such as citation, webpage, and product graphs feng2024taglas ; li2025hetgb , have nodes associated with text attributes. Graph neural networks (GNNs) kipf2016semi ; yun2019graph have demonstrated excellent performance and efficiency in node classification on TAGs, which are supported by high-quality labels and effective message-passing mechanisms zhu2021textgnn . However, the manual labeling of nodes is undoubtedly a tedious, expensive, and time-consuming task shi2025latent ; cao2025analytical . In many scenarios, only a few node labels are available. Additionally, nodes often have complex and diverse interaction relationships with each other—their local topologies are intricate—which challenge traditional GNNs with fixed message-passing mechanisms. Compared with GNNs, large language models (LLMs) exhibit impressive zero-/few-shot learning capabilities on TAGs chen2024exploring ; wang2024llms ; wu2025exploring . But the large parameter scale considerably hinders their inference efficiency he2024harnessing .
A natural idea is to blend their complementary strengths for few-shot node classification on TAGs. Knowledge distillation (KD) hinton2014distilling is a feasible solution. However, directly distilling knowledge from the LLM to GNN is impractical. Firstly, the discrepancy of decoder-only (LLMs) and encoder-only (GNNs) leads to fundamentally different characteristics in their embedding spaces chenlabel . And the huge embedding-dimension difference needs sophisticated embedding alignment and also brings high training cost hu2025large . In contrast, conducting prediction distillation from LLMs to GNNs by annotating node labels can efficiently alleviate the label scarcity and scalability dilemma pan2024distilling . The critical question is how to select the nodes for the LLM’s label annotation to effectively enhance teacher GNNs. Generally, one may use uncertainty fox1957training as a selection metric in the embedding space of GNN. However, owing to nodes’ diverse semantic and complex structural attributes (e.g., local topologies), a single GNN cannot capture the essences of nodes completely liu5084903distilling . Therefore, we investigate the embedding spaces of various-architecture GNNs to effectively mitigate cognitive limitations wang2021mulde associated with relying on a single GNN, thereby better selecting nodes for LLM’s label annotations.
@@
To this end, we propose a preference-driven knowledge distillation (PKD) framework that unites the complementary strengths of LLMs and various-architecture GNNs for few-shot node classification on TAGs. It mainly includes two modules: GNN-preference-driven Node Selector (GNS) and Node-preference-driven GNN Selector (NGS). The prerequisite of GNS is that the LLM should be able to comprehend the graph topology. Thus, we develop the graph topology aware (GTA) prompts to fine-tune the LLM, enhancing its capacity to comprehend graph topology. GNS fully exploits nodes’ prediction discrepancies among various GNNs to decide nodes whose labels are annotated by the LLM will effectively enhance teacher GNNs, facilitating knowledge distillation from the LLM to teacher GNNs. NGS selects for each node the most appropriate GNN message-passing mechanism, facilitating the tailored knowledge distillation from various teacher GNNs to the student GNN. It regards the fine-tuned LLM as the RL-based (reinforcement learning) agent, which treats all textualized node-specific attributes (including node’s semantic, structure, and prediction attributes) as state and the student GNN’s performance as reward. Our contributions can be summarized as follows:
  • We introduce a preference-driven knowledge distillation (PKD) framework to synergize the complementary strengths of the LLM and various GNNs ingeniously for few-shot node classification on TAGs;
  • We propose a GNN-preference-driven node selector, effectively determining nodes for annotation by the LLM and promoting knowledge distillation from the LLM to teacher GNNs;
  • We propose a node-preference-driven GNN selector to tailor for each node the most appropriate message-passing mechanism, promoting knowledge distillation from teacher GNNs to the student GNN;
  • We validate the efficacy of PKD for few-shot node classification on nine TAGs. The experiments show that it even defeats some state-of-the-art methods that use more node labels.
@@ 2.1 Graph Neural Networks
The field of graph learning has been dominated by GNNs. Early GCN kipf2016semi introduces a spectral-based graph convolution operation to propagate node information through the graph. GAT velivckovic2018graph uses attention mechanisms to weigh neighbors’ contributions, enabling adaptive learning of neighborhood importance. APPNP gasteiger2018combining enhances message passing by using personalized propagation with a power iteration approach, improving label propagation on graphs. $\text{H}_{2}\text{GCN}$ zhu2020beyond extends GCNs by incorporating higher-order neighborhood information to improve representation power. GPRGNN chien2021adaptive combines graph convolution with residual connections to improve propagation efficiency, particularly in graphs with diverse node degrees. HoloNets koke2024holonets introduces a dual-filter mechanism with spectral response, extending spectral convolutions to directed graphs. DirGNN rossi2024edge defines the in-neighbors and out-neighbors and performs separate propagation and aggregation, improving the message passing through the incorporation of edge directionality. To deal with label scarcity, GCNII chen2020simple introduces initial residual connections and identity mapping to construct a deep GNN while EGNN satorras2021n enforces equivariance constraints for the enhancement of data efficiency and generalization. AGST ding2024toward and IceBerg li2025iceberg leverage the different self-training li2018deeper methods to effectively utilize unlabeled nodes.
@@ 2.2 Knowledge Distillation
KD is not only used for model compression, but for strengthening purposeful abilities of the student model. GFL yao2020graph extracts structural knowledge from a pre-prepared similar auxiliary graph, distilling it to the target graph for enhancing few-shot node classification performance. KDGA wu2022knowledge utilizes multiple graph augmentation strategies to make student GNN produce robust node representations after distillation. MSKD zhang2022multi mitigates the diverse classification situations requiring for different nodes by capturing multi-scale topological semantics distilled from varying layers. However, the capability of an individual teacher is inherently limited. BGNN guo2023boosting distills complementary knowledge from multiple GNN teachers sequentially and integrate it by the adaptive temperature parameter and weight boosting modules. MTAAM yang2024two distills knowledge of multiple teacher GNNs into an MLP-student, offering quick inference speed without compromising accuracy. FairGKD zhu2024devil obtains equitable and informative node representations by synergizing multiple GNN experts into a teacher. DMKD liu5084903distilling harnesses complementary knowledge from various GNNs and conducts layer-level knowledge distillation to mitigate the constraint of a single teacher. Furthermore, chenlabel is a label-free method that proposes the LLM-GNN. It uses LLMs to get high-quality annotation through active and confidence-awareness node selection, thereby circumventing the difficulty of label annotation by humans. LinguGKD hu2025large introduces a kind of ingenious contrastive learning to align the LLM’s semantic features with GNN’s structural features to achieve knowledge transfer. Most of the above knowledge distillation methods do not tailor for each node the most appropriate message-passing mechanism and underperform on few-shot node classification.
@@ 4.1 Experimental Setup
In order to assess the few-shot node classification performance of our method on TAGs, we conduct a comprehensive series of experiments across 9 real-world datasets: Cornell, Washington, Texas, Wisconsin zhu2020beyond , Amazon Ratings platonov2023a , Ogbn-Arxiv hu2020open , Wiki CS mernyei2020wiki , Pubmed, Cora yang2016revisiting . They have various 1-hop homophily ratios yang2024incorporating and additional details of the datasets can be found in Appendix A. For the KD-baselines, we partition the nodes of each graph into training, validation, and test sets, allocating 48%, 32%, and 20%, respectively, based on the proportion division mentioned in pmlr-v235-liang24c . For PKD and other baselines, we randomly select 1, 3, and 5 labeled nodes per class as the initial training data and then expand the dataset to 48% of the total using the GNS module. The remaining data is randomly split into 32% for validation and 20% for testing, with the preserved indices for the baselines. This operation is repeated 5 times. We report the average test classification accuracy and standard deviation of each model with parameters that lead to the peak validation accuracy.
@@
We compare our method against the following baseline models: (i) Advanced GNNs: GCNII chen2020simple and EGNN satorras2021n ; (ii) GNNs enhanced by LLMs: LLMGNN chenlabel and GAugLLM fang2024gaugllm ; (iii) self-training for graph learning: Self-training li2018deeper , AGST ding2024toward and IceBerg li2025iceberg ; (iv) Knowledge Distillation (KD) for GNNs: KDGA wu2022knowledge , MSKD zhang2022multi , BGNN guo2023boosting , MTAAM yang2024two , and FairGKD zhu2024devil . For homophily graphs, the teacher GNNs used are: GCN kipf2016semi ($T_{1}$), GAT velivckovic2018graph ($T_{2}$), APPNP gasteiger2018combining ($T_{3}$), $\text{H}_{2}\text{GCN}$ zhu2020beyond ($T_{4}$), and the student is GCN; for heterophily graphs, the teacher GNNs employed are: DirGNN rossi2024edge ($T_{1}$), GPRGNN chien2021adaptive ($T_{2}$), HoloNets koke2024holonets ($T_{3}$), $\text{H}_{2}\text{GCN}$ ($T_{4}$) and the student is $\text{H}_{2}\text{GCN}$. The LLM used in the experiments is Llama-3.1-8B-Instruct dubey2024llama .
@@ 45.02 Amazon Ratings
To assess the effectiveness of each part in the reward function (Eqn. (4)), we visualize the training processes of three variants in Figure 4: (a) $R_{1}$: The reward function for teacher GNN selection depends solely on the classification accuracy of the student GNN on the expanded training data; (b) $R_{2}$: In addition to classification accuracy, the reward function also incorporates the negative cross-entropy loss ($\mathcal{L}_{CE}$); (c) $R_{3}$: Building upon $R_{2}$, the reward function also includes the negative knowledge distillation loss ($\mathcal{L}_{DL}$). As shown in Figure 4, both the three parts contribute to the improved classification performance.
@@ 4.5 Running Time
Datasets / Methods Self-training AGST IceBerg KDGA MSKD BGNN MTAAM FairGKD PKDLlama
@@ Appendix A Detailed Description of Datasets
Cornell, Washington, Texas, and Wisconsin:
@@
Cora, Pubmed, and Ogbn-Arxiv:
@@ B.1 Prompts for Zero-shot Node Classification
Table 9: The prompt template for zero-shot node classification.
@@ B.2 Prompts for Node-preference-driven GNN Selector
Table 10: The prompt template for Node-preference-driven GNN Selector.
@@ B.3 Graph Topology Aware (GTA) Prompts
The full prompts for Connectivity is presented above. When generating prompts for different datasets, we adjust certain descriptions to better align with the specific context. For example, when constructing prompts for Texas, the background description should be adapted to reflect web pages, and the relationship should be revised to hyperlinks, along with other context-specific adjustments. Similarly, for each task, the prompts must also be modified to correspond to the specific content described in Section 3.2.
@@ Appendix C Implementation Details and Time Complexity Analysis
1 $\tilde{\mathcal{D}_{L}}\leftarrow\mathcal{D}_{L}$;
@@ 11      Store ($\mathcal{P}_{i},\mathbf{m}_{i},R_{i}$) to the episode history $\mathcal{F}$;
The time complexity of PKD training is mainly divided into three parts: LLM fine-tuning (Line 2), GNN-preference-driven Node Selector (Line 3-4) and Node-preference-driven GNN Selector (Line 5-12). The GNN-preference-driven Node Selector also can be divided into the annotations generation and prediction distillation.
@@
Model Name Dataset Size Epoch lora_r lora_alpha Optimizer Learning Rate Time Cost
Llama-3.1-8B-Instruct 53,617 2 4 4 AdamW [57] $1e^{-4}$ 9h 41m 48s
@@
Table 13: The time costs on Cora and Ogbn-Arxiv of annotations generation by Llama-3.1-8B-Instruct.
Dataset Cora Ogbn-Arxiv
Time / GPU-hours 0.11 1.68
@@
Table 14: The time costs of retraining teacher GNNs on Cora and Ogbn-Arxiv.
Datasets Teacher GNNs ($T_{1},T_{2},T_{3},T_{4}$) Total running time (seconds)
Cora GCN, GAT, APPNP, H2GCN 2.4716
Obgn-Arxiv GCN, GAT, APPNP, H2GCN 18.2017
@@
Table 15: Running times of Node-preference-driven GNN Selector on Cora and Obgn-Arixv. "m" and "s" denote minute and second.
Dataset Peak memory Running time / epoch
Cora 454.62 MB 7.9s
Obgn-Arixv 1655.18 MB 44m 8.3s
The time complexity of Node-preference-driven GNN Selector is $\mathcal{O}(W(l^{2}d+ld^{2}+dd’+d^{\prime}a))$, where $d^{\prime}$ is the dimension of the MLP hidden layer, $a$ is the number of action categories, $W\ll l$, $a\ll d$, so the time complexity is bound by $\mathcal{O}(l^{2}d+ld^{2}+dd^{\prime})$. The training time complexity of student GNN is $\mathcal{O}((NF+M)D)$. Therefore, the overall time complexity is bound by $\mathcal{O}((L+2l)d^{2}+(d’+nL)d+2l^{2}+2D(NF+M))$.